OpenStudy (anonymous):

Frank, carl, and vinny started a lawn mowing business over the summer. This table shows how many lawns each worker mowed during summer months. | worker | June | July | Aug | | frank...| 28....| 21....| 19....| | carl.....| 17....| 24....| 15....| | vinny...| 22....| 29....| 17....| what is element w32 A. 19 B.29 C.22 D.21

OpenStudy (anonymous):

what is w32?

OpenStudy (anonymous):

it just says display the data in matrix W with rows indicating workers.

OpenStudy (anonymous):

ummm not to sure on this one

OpenStudy (anonymous):

actually what we have to ultimately find?

OpenStudy (anonymous):

W32

OpenStudy (anonymous):

29

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[\left[\begin{matrix}28 & 21&19 \\ 17 & 24& 15\\22&29&17\end{matrix}\right]\]

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

yeah i think so. thank you again!

OpenStudy (anonymous):

i'll give the reference\[\left[\begin{matrix}w11 & w12&w13 \\ w21 & w22&w23\\w31&w32&w33\end{matrix}\right]\]

OpenStudy (anonymous):

is this ok?

OpenStudy (anonymous):

yes!